3. The demand for cookies is given by P = 36.2 – 0.2QD where QD is the quantity demanded of the cookies (number of cookies per day) and P is the price of cookies ($/cookie). Also, the supply of cookies is given by P = 1.4 + 0.3QS where QS is the quantity supplied of the cookies (number of cookies per day) and P is the price of cookies ($/cookie).

Find the equilibrium price and quantity. Show your work.
1
Expert's answer
2020-09-21T09:06:26-0400

"At Equilibrium,\nQD = QS"

"36.2 - 0.2Q = 1.4 + 0.3Q"

"36.2 - 1.4 = 0.3Q + 0.2Q"

"34.8 = 0.5Q"

"Q = 34.8\/0.5 = 69.9"

"P = 36.2 - 0.2 (69.9) = 22.28"


Therefore, Equilibrium Quantity (EQ) is 69.9 cookies per day, while Equilibrium Price (EP) $22.28/cookie